OpenStudy (anonymous):

Completing the Square of x²=2x+33 ?

OpenStudy (anonymous):

x^2 = 2 x+33 Subtract (2 x+33) from both sides: x^2-2 x-33 = 0 Add 33 to both sides: x^2-2 x = 33 Add 1 to both sides: x^2-2 x+1 = 34 Factor the left hand side: (x-1)^2 = 34 Take the square root of both sides: sqrt(x-1) = sqrt(34) Eliminate the absolute value: x-1 = -sqrt(34) or x-1 = sqrt(34) Add 1 to both sides: x = 1-sqrt(34) or x-1 = sqrt(34) Add 1 to both sides: x = 1-sqrt(34) or x = 1+sqrt(34)

OpenStudy (earthcitizen):

Dashing!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so x = 1-sqrt(34) or x = 1+sqrt(34) is the final answer ?

OpenStudy (anonymous):

yep